A negative feedback system has a forward transfer function 1/(s+7.8) and feedback transfer function s+1.8. Obtain the time constant (in sec) of the overall transfer function.

Give the answer in 4 decimal places. (Unit is not required.)
Answer #1

→ Giren GCS) = H(s) = 5+1.8 4 H(s) = (st $+7.8) → The closed loop transfer function is GC) = 0.104 I +G(s) H(s) 25+ 9.6 (+ 25